Dyrdek Enterprises has equity with a market value of $11.8 million and the market value of debt is $4.05 million. The company is evaluating a new project that has more risk than the firm. As a result, the company will apply a risk adjustment factor of 2.1 percent. The new project will cost $2.40 million today and provide annual cash flows of $626,000 for the next 6 years. The company's cost of equity is 11.47 percent and the pretax cost of debt is 4.98 percent. The tax rate is 35 percent. What is the project's NPV?

$377,779

$213,119

$559,481

$212,843

$180,192
Answer #1

Total market value = 11.8 + 4.05 = 15.85

WACC = Weight of debt*after tax cost of debt + weight of equity*cost of equity

WACC = (4.05 / 15.85)*0.0498*(1 - 0.35) + (11.8 / 15.85)*0.1147

WACC = 0.008271 + 0.085392

WACC = 0.093663 or 9.3663%

WACC to be used = 9.3663% + 2.1% = 11.4663%

NPV = Annuity * [1 - 1 / (1 + r)n] / r - Initial investment

NPV = 626,000 * [1 - 1 / (1 + 0.114663)6] / 0.114663 - 2,400,000

NPV = 626,000 * 4.174311 - 2,400,000

NPV = $213,119